A farmer mixes a fertilizer solution that uses 15 parts water to 5 parts fertilizer. If he mixes a solution that uses 10 gallons of fertilizer, how many gallons of water will he use?

Answer by checkley79(3341)

15/5=X/10 CROSS MULTIPLY.
5X=15*10
5X=150
X=150/5
X=30 GALLONS OF WATER ARE REQUIRED.
PROOF:
15/5=30/10
3=3